RC Update and Run Plan for Aug 10, 2020
- 1 RC Meeting Information
- 2 General updates
- 3 Current Run Plan
- 4 Notices in Effect for Shift Crew
- 5 Planned Accelerator Configuration Changes
- 6 Useful Links
RC Meeting Information
RC Daily Meetings are at 4:15 PM, remote
Monday Aug 10 2020.
- Finally got beam at the end of swing. Able to do beam checkout. Very trippy.
- Requested 27 uA of beam through the night and made changes as required to improve stability.
- Beam seemed to improve on day shift.
- Hall D had a down, so we took BCM calibration. Highest current 34 uA. 
- HMS changed yesterday only in epics readback from 20.0 to 20.03--weird timing 
Recent polarization measurements
- 42.8% at 3:15pm 
Current Run Plan
Normal Production running
- Make sure operational restrictions are followed.
- Make sure the raster size is correct (4-4.5)
- Monitor beam parameters/configuration very often and post an entry like this at the beginning of your shift and when run conditions have changed
- Running on cell targets: raster always ON (4-4.5, size will be determined by the scraping checks), orbit locks ON, current ramp rate 1 muA/sec
- Always take 120Hz of EDTM
- Fill the BTA with realistic numbers. If the desired beam current is not delivered prorate your ABUs!
- Change the IHWP every 4 hours when running on glass cells
- Run plan
- Production on 3He; if not able to reach requested current (27 uA), pro-rate the ABUs
Notes for tomorrow's 7:45 meeting
- Opportunistic target polarization calibration and studies on Tuesday-Wednesday during beam off and Moller measurement (priority is beam)??
- Moller measurement on Wednesday
- Plan for BCM calibration when C has priority. Need two more cycles (disruptive to D due to shared slit).
- After ~ 10 days of running Hall C will go to 1 pass for ~3 days
- EPR diode good for now, but check for future planned accesses.
- Plan for Tuesday--Controlled Access:
- Beam ends.
- Target to change spin polarization (270->90). EPR calibration.
- RADCON sweeps Hall at 8am. Steve L enables rotation and enters with Radcon to rotate HMS from 20 to 29 deg.
- Simona, Walter and techs to enter Hall for SoLID cerenkov and HMS Q1 vacuum check. Anticipated 4-5 hours of work.
- Steve L. will remotely update the software on HMS dipole to decrease interlock due to false spike in current lead mass flow .
When there is a down:
Notices in Effect for Shift Crew
- Beam positions: 
- Policy for CANS-related hall status change: Make a logbook entry when the hall goes in, or comes out of laser-controlled access.
- Magnet policies
- Spectrometer GUI (
go_magnets) should always be visible on Shift Leader computer, hcdesk3
- HMS Q1 must be ramped down in 50A steps in all cases
- Do not ramp HMS Q2 to zero, only go to 10A and then switch off power.
- If any of the Spectrometers magnets trip, take a screenshot of the PSU window and check the hall for smoke or water with the camera before resetting the interlock and post it to hclog.
- Spectrometer GUI (
- Inform the RC of any access that the Target Team requires. The RC will coordinate with RadCon to make everything run smoothly.
- Please make sure that the orbit locks are enabled and the beam does not drift from its nominal position.
- Ramp rate lock instructions are here and added to checklist
Planned Accelerator Configuration Changes
- Bluejeans remote connection information:
To join the Meeting: https://bluejeans.com/861439920 To join via Room System: Video Conferencing System: bjn.vc -or-220.127.116.11 Meeting ID : 861439920 To join via phone : 1) Dial: +1.408.740.7256 (US (San Jose)) +1.888.240.2560 (US Toll Free)) +1.408.317.9253 (US (Primary, San Jose)) (see all numbers - http://bluejeans.com/numbers) 2) Enter Conference ID : 861439920